# Integrated workflows and interfaces for data-driven semi-empirical electronic structure calculations

### v1

@@ Abstract
Modern software engineering of electronic structure codes has seen a paradigm shift from monolithic workflows towards object-based modularity. Software objectivity allows for greater flexibility in the application of electronic structure calculations, with particular benefits when integrated with approaches for data-driven analysis. Here, we discuss different approaches to create “deep” modular interfaces that connect big-data workflows and electronic structure codes, and explore the diversity of use cases that they can enable. We present two such interface approaches for the semi-empirical electronic structure package, DFTB+. In one case, DFTB+ is applied as a library and provides data to an external workflow; and in another, DFTB+ receives data via external bindings and processes the information subsequently within an internal workflow. We provide a general framework to enable data exchange workflows for embedding new machine-learning-based Hamiltonians within DFTB+, or to enabling deep integration of DFTB+ in multiscale embedding workflows. These modular interfaces demonstrate opportunities in emergent software and workflows to accelerate scientific discovery by harnessing existing software capabilities.
@@ I Introduction
The majority of well-established electronic structure software packages have developed as monolithic codebases with a single entry point, due to extensive investment in their development before the widespread adoption of integrated workflows. The consequence is a bottleneck for the integration between electronic structure and big-data approaches. Electronic structure software packages with modularity in their design, that provide external accessibility to inner functionality, have become more prominent in recent years, reflecting the evolving landscape of computational materials science and the uptake of objective, modular programming and data-driven workflows (e.g., GPAW,[23] Psi4,[24] and pySCF[25]) or UNIX philosophy[26] inspired designs such as WIEN2k[27]; furthermore, established packages have sought to reshape their designs with library components that allow execution through an externally-driven interface. These retrofitted approaches typically rely on file input/output (I/O) and parsing, with only basic variable communication (e.g., MPI communicators), though alternative strategies with in-memory data transmission have increased in popularity. Examples of established strategies include the automated building of deep Python interfaces to Fortran codes using f90wrap[28] and socket-type interfaces.[29, 30] High-level Pythonic packages, such as the Atomic Simulation Environment (ASE)[31], have emerged as a de facto standard for building atomistic simulation workflows,[32], enabling some classes of high-level algorithms to be written once and reused between codes; however, this generality is commonly restricted to atomic and molecular properties rather than electronic. Recently, the emergence of automated and machine-learning-augmented workflows, and establishment of extensive materials databases using FAIR principles,[33] has led to a need for more flexible infrastructure capabilities in the design of electronic structure software. In particular, there is evidence of a clear need for greater modularity and interoperability in code design, which should support strong interfaces between electronic structure codes and external software packages. Such developments would circumvent traditional bottlenecks in data communication and accelerate discoveries facilitated by electronic structure theory.
@@
Currently, there remains limited demonstration and standardization of deployable, user-ready interfaces that facilitate interaction and application of external workflows and data-driven frameworks with first-principles and semi-empirical electronic structure software packages. In particular, there is a need for “deep” module interfaces that can expose the extensive and well-developed functionality in existing established software. The interfaces should be simple code-level interfaces, rather than commonly shallow interfaces that work predominantly with complex file I/O or external scripted workflows. The ability to use such “deep” interfaces, which allows large data objects to be manipulated during run-time, will reward the community with computationally efficient approaches concerning both data processing and data storage. The Atomic Simulation Interface (ASI)[34] is a recent example interface that is built to import and export electronic structure information from quantum chemistry codes during runtime with minimal performance penalty, as demonstrated via coupling with the DFTB+ and FHI-aims software packages.[35]
@@
Recent work on the DFTB+ software package[36, 37] has investigated how “deep” interfaces can be established and used for the benefit of workflow-based computational simulation. Herein, we discuss general strategies for interfacing to electronic structure codes before presenting two interfaces that are capable of either being driven by an external workflow to provide data,[34] or driving a workflow, with external bindings used to obtain data from an external engine.[18] The interfaces follow different strategies and philosophies, and address distinct potential use cases, such as embedding and ML workflows.
@@ II General considerations on deep interfaces to electronic structure codes
Approaches to suspend and restart a subroutine map on to two different code interfacing paradigms (see Figure 1). One option is refactoring and splitting code into two separate subroutines that are called immediately before and after any data import or export. An alternative option is invoking a user-provided callback function to perform data read or write. The former method essentially converts the code into a library, inverting the control of the workflow; if the routine that is split is nested deep within the call stack, every function along the call stack must also be split. The latter method introduces local inversions of control through callbacks, but the modified code generally still drives the overall workflow. In the following, we present realizations of these paradigms inside the DFTB+ software package.
@@ V Atomic Simulation Interface (ASI) as a driver that uses DFTB+
The ability to drive calculations externally, and use a specific package to evaluate system properties on demand, motivates the development of an infrastructure where DFTB+ can be deployed as a software library. Modern Python coding developments provide capacity for high-level interfaces, reliant on file I/O for data transfer, but “deep” integration via pre-compiled software languages can enable more efficient and accurate data communication and software application. Recent efforts towards this software paradigm have seen the development of the Atomic Simulation Interface (ASI), with the primary purpose of conveniently connecting ASI-enabled codes in multiscale simulation workflows, such as hybrid quantum/molecular mechanics (QM/MM), multiscale quantum mechanical embedding (QM/QM), or integration with machine learning (ML) frameworks (QM/ML).
@@ V.1 Interface Description
The key ASI functions can be broken into four groups: control flow; atomic information; electrostatic potential; and electronic structure matrices.[34] The necessary intrusions to implement in an existing codebase are minimal. For the application of DFTB+ using the ASI standard, the DFTB+ package is compiled as a shared object library to allow dynamic linkage with the client. The workflow is driven by the ASI client; thus, after ASI initialisation, key data objects are communicated to/from DFTB+ and callback functions registered before the request for execution of a DFTB+ calculation. Callback functions give direct access to data objects within DFTB+, thus causing near-zero computational cost, and also adapting to the chosen parallelisation schemes. The callback functions are invoked during the execution process, providing external access to data objects when calculated. Derived quantities, such as energy, forces, stress, atomic charges, are also available. Once all necessary operations on the exposed data objects have been completed, finalization is performed, which includes the release of allocated memory. The ASI workflow is presented in Figure 2, and contrasted against the ACEhamiltonians interface. Further details of the key ASI functions are provided in the SI.
@@ VI Conclusions
The DFTB+ software package is available at https://github.com/dftbplus/dftbplus. The v24.2 release will contain all functionality outlined in this manuscript; the described changes for the ASI binding or to connect to the ACEhamiltonians are undergoing review and are currently available a [61] and [62] respectively. Full documentation is available at https://dftbplus.org/. The ACEhamiltonians v0.1.0 software package is available at https://github.com/ACEsuit/ACEhamiltonians.jl. The ASI v1.1 software package is available at https://gitlab.com/pvst/asi. The interface specification and DFTB+ implementation are available at https://pvst.gitlab.io/.

### v2

@@ Abstract
Modern software engineering of electronic structure codes has seen a paradigm shift from monolithic workflows towards object-based modularity. Software objectivity allows for greater flexibility in the application of electronic structure calculations, with particular benefits when integrated with approaches for data-driven analysis. Here, we discuss different approaches to create deep modular interfaces that connect big-data workflows and electronic structure codes, and explore the diversity of use cases that they can enable. We present two such interface approaches for the semi-empirical electronic structure package, DFTB+. In one case, DFTB+ is applied as a library and provides data to an external workflow; and in another, DFTB+ receives data via external bindings and processes the information subsequently within an internal workflow. We provide a general framework to enable data exchange workflows for embedding new machine-learning-based Hamiltonians within DFTB+, or to enabling deep integration of DFTB+ in multiscale embedding workflows. These modular interfaces demonstrate opportunities in emergent software and workflows to accelerate scientific discovery by harnessing existing software capabilities.
Statement of Significance
Software is integral to modern science, and modularity and interoperability of software will allow new capabilities that are greater than the sum of the constituent elements. We demonstrate such capabilities via the integration of modular software packages with the semi-empirical electronic structure package DFTB+. Our work is exemplar in showing how a well-developed software package, such as DFTB+, can undergo re-engineering to allow work package and workflow interoperability in either client or server capacity. It is imperative that the community of chemical physics embraces the opportunities presented by modular software design and interoperability of software packages; our work is therefore significant because it provides original demonstrations of how modular software integration can lead to higher-value outcomes from electronic structure calculations. These new capabilities can be used to guide further developments in this field and related applications.
@@ I Introduction
The majority of well-established electronic structure software packages have developed as monolithic codebases with a single entry point, due to extensive investment in their development before the widespread adoption of integrated workflows. The consequence is a bottleneck for the integration between electronic structure and big-data approaches. Electronic structure software packages with modularity in their design, that provide external accessibility to inner functionality, have become more prominent in recent years, reflecting the evolving landscape of computational materials science and the uptake of objective, modular programming and data-driven workflows (e.g., GPAW,[23] Psi4,[24] and pySCF[25]) or UNIX philosophy[26] inspired designs such as WIEN2k[27]; furthermore, established packages have sought to reshape their designs with library components that allow execution through an externally-driven interface. These retrofitted approaches typically rely on file input/output (I/O) and parsing, with only basic variable communication (e.g., MPI communicators), though alternative strategies with in-memory data transmission have increased in popularity.
Examples of established strategies include the automated building of deep Python interfaces to Fortran codes using f90wrap[28] (e.g., QEPy for Quantum Espresso [29], quippy for QUIP [30], CasPyTep for CASTEP[31], and a Python wrapper for the Bader code [32]), socket-type interfaces (i-PI [33], MDI [34]), and library extraction (tblite in CREST [35], ELSI [36]). High-level Python packages, like the Atomic Simulation Environment (ASE)[37], have emerged as a de facto standard for building frameworks for atomistic simulation workflows. Examples of workflow engines include AiiDA [38], Chemoton [39], CENSO [40], BigChem [41], QCEngine [34], etc., which enable some classes of high-level algorithms to be written once and reused between codes; however, this generality is commonly restricted to atomic and molecular properties rather than electronic structure.
Recently, the emergence of automated and machine-learning-augmented workflows, and establishment of extensive materials databases using FAIR principles,[42] has led to a need for more flexible infrastructure capabilities in the design of electronic structure software. In particular, there is evidence of a clear need for greater modularity and interoperability in code design, which should support strong interfaces between electronic structure codes and external software packages. Such developments would circumvent traditional bottlenecks in data communication and accelerate discoveries facilitated by electronic structure theory.
@@
Currently, there remains limited demonstration and standardization of deployable, user-ready interfaces that facilitate interaction and application of external workflows and data-driven frameworks with first-principles and semi-empirical electronic structure software packages. In particular, there is a need for deep module interfaces that can expose the extensive and well-developed functionality in existing established software. The interfaces should be simple code-level interfaces, rather than commonly shallow interfaces that work predominantly with complex file I/O or external scripted workflows. The ability to use such deep interfaces, which allows large data objects to be manipulated during run-time, will reward the community with computationally efficient approaches concerning both data processing and data storage. The list of software that may benefit from using deep interfaces includes charge partitioning codes [43, 44, 45], electronic structure analysis [46], and the growing family of electronic structure machine learning models [18, 22, 20, 19]. The Atomic Simulation Interface (ASI)[47] is a recent example interface that is built to import and export electronic structure information from quantum chemistry codes during runtime with minimal performance penalty, as demonstrated via coupling with the DFTB+ and FHI-aims software packages.[48]
@@
Despite the recognised disadvantages of file I/O-based interfacing, this approach is the most commonly used option. Reasons that file I/O interfacing remains popular is that the file system interface is universally provided by the operating system, it does not impose any restrictions on data format, and it is intrinsically asynchronous. These features are useful if codes on both ends of the interface are not specified a priori, which is especially applicable in the case of closed-source software. Motivations for deeper interfaces include the complexity of I/O data formats and the stability of their layout over time, the synchronisation of I/O data, and inherent performance limitations when using I/O, especially at high frequency. The ongoing efforts towards standardisation of input and output formats, such as QCEngine [34], JSON [49] in ORCA [50] or HDF5-based TREXIO [51], offer potential to alleviate some of these challenges.
Recent work on the DFTB+ software package[52, 53] has investigated how deep interfaces can be established and used for the benefit of workflow-based computational simulation. Herein, we discuss general strategies for interfacing to electronic structure codes before presenting two interfaces that are capable of either being driven by an external workflow to provide data,[47] or driving a workflow, with external bindings used to obtain data from an external engine.[18] The interfaces follow different strategies and philosophies, and address distinct potential use cases, such as embedding and ML workflows.
@@ II General considerations on deep interfaces to electronic structure codes
Approaches to suspend and restart a subroutine map on to two different code interfacing paradigms (see Figure 1). One option is refactoring and splitting code into two separate subroutines that are called immediately before and after any data import or export. An alternative option is invoking a user-provided callback function to perform data read or write. The former method essentially converts the code into a library, inverting the control of the workflow; if the routine that is split is nested deep within the call stack, every function along the call stack must also be split. The latter method introduces local inversions of control through callbacks, but the modified code generally still drives the overall workflow.
In general, the considerations outlined may be viewed as significant drawbacks to deep interfaces, especially if the interfaced code is poorly structured, fragile, or tightly coupled; in such cases, development and maintenance of deep interfaces may become involved and laborious. However, the potential benefits are significant and, if pursued carefully, present opportunities for realising new science. In the following, we present realisations of the two highlighted deep interfacing paradigms inside the DFTB+ software package, outlining our experiences with realisation and showcasing the potential value.
@@ V Atomic Simulation Interface (ASI) as a driver that uses DFTB+
The ability to drive calculations externally, and use a specific package to evaluate system properties on demand, motivates the development of an infrastructure where DFTB+ can be deployed as a software library. Modern Python coding developments provide capacity for high-level interfaces, reliant on file I/O for data transfer, but deep integration via pre-compiled software languages can enable more efficient and accurate data communication and software application. Recent efforts towards this software paradigm have seen the development of the Atomic Simulation Interface (ASI), with the primary purpose of conveniently connecting ASI-enabled codes in multiscale simulation workflows, such as hybrid quantum/molecular mechanics (QM/MM), multiscale quantum mechanical embedding (QM/QM), or integration with machine learning (ML) frameworks (QM/ML).
@@ V.1 Interface Description
The key ASI functions can be broken into four groups: control flow; atomic information; electrostatic potential; and electronic structure matrices.[47] The necessary intrusions to implement in an existing codebase are minimal. For the application of DFTB+ using the ASI standard, the DFTB+ package is compiled as a shared object library to allow dynamic linkage with the client. The workflow is driven by the ASI client; thus, after ASI initialisation, key data objects are communicated to/from DFTB+ and callback functions registered before the request for execution of a DFTB+ calculation. Callback functions give direct access to internal data objects within DFTB+ process via pointers, without unnecessary copying, thus causing near-zero computational and communication overhead, and also adapting to the chosen parallelisation scheme. The callback functions are invoked during the execution process, providing external access to data objects when calculated. Derived quantities, such as energy, forces, stress, atomic charges, are also available. Once all necessary operations on the exposed data objects have been completed, finalization is performed, which includes the release of allocated memory. The ASI workflow is presented in Figure 2, and contrasted against the ACEhamiltonians interface. Further details of the key ASI functions are provided in the SI.
@@ VI Conclusions
Supplementary Material
The accompanying supplementary information provides: a comparison of band structures calculated with FHI-aims and DFTB+; details of the DFTB+ external API; details of the ASI API.
@@
The DFTB+ software package is available at https://github.com/dftbplus/dftbplus. The v24.2 release will contain all functionality outlined in this manuscript; the described changes for the ASI binding or to connect to the ACEhamiltonians are undergoing review and are currently available at [76] and [77] respectively. Full documentation is available at https://dftbplus.org/. The ACEhamiltonians v0.1.0 software package is available at https://github.com/ACEsuit/ACEhamiltonians.jl. The ASI v1.1 software package and DFTB+ implementation are available at https://gitlab.com/pvst/asi. The interface specification is available at https://pvst.gitlab.io/asi.